### 30 Mar 2026 v1

# ITQ3_S: Interleaved Ternary Quantization with TurboQuant High-Fidelity 3-bit LLM Inference via Rotation-Domain Adaptive Quantization

Abstract
We present ITQ3_S (Interleaved Ternary Quantization – Specialized), a novel 3-bit weight quantization format for large language models (LLMs) that integrates TurboQuant (TQ), a rotation-domain adaptive quantization strategy based on the Fast Walsh-Hadamard Transform (FWHT). Conventional 3-bit quantization methods suffer from catastrophic precision loss caused by heavy-tailed weight distributions and inter-channel outliers. ITQ3_S addresses this fundamental limitation by pre-rotating the weight space via FWHT prior to quantization, effectively spreading outlier energy across the entire vector and inducing a near-Gaussian distribution amenable to uniform ternary coding.
Critically, we derive a mathematically rigorous dequantization procedure that inverts the FWHT exactly using a 256-point Inverse Walsh-Hadamard Transform fused into the CUDA shared-memory loading stage, ensuring zero-error round-trip fidelity between offline quantization and online inference. We prove that for any weight vector $\mathbf{w}\in\mathbb{R}^{256}$ processed by our pipeline, the reconstruction satisfies $\|\hat{\mathbf{w}}-\mathbf{w}\|_{2}\leq\epsilon_{q}$, where $\epsilon_{q}$ is determined solely by the ternary quantization grid and is strictly smaller than any uniform 3-bit baseline under equal bit-budget constraints.
[…]
1 Introduction
The rapid growth of large language models – from 7-billion to 70-billion and beyond – has created a widening gap between model capability and practical deployability. While state-of-the-art LLMs exhibit remarkable reasoning abilities, their memory footprint in FP16 or BF16 precision exceeds what is feasible on even the most powerful consumer GPUs. A 70B-parameter model in FP16 requires approximately 140 GiB of memory; even the NVIDIA RTX 5090 with 32 GiB of VRAM cannot load it without quantization.
[…]
The transition from 4-bit to 3-bit quantization is especially challenging. Empirically, 3-bit has been called the “breaking point” for LLM logic: models quantized naively to 3 bits exhibit sharp perplexity degradation, hallucination, and loss of multi-step reasoning capability. This degradation stems from two sources:
  1. 1.
    Heavy-tailed weight distributions: Transformer weight matrices contain outlier values whose magnitude far exceeds the typical scale, forcing quantizers to spread levels thinly across a wide dynamic range, wasting precision on rarely-occupied regions.
  2. 2.
    Inter-channel correlation: Structured correlation among weight channels causes uniform quantization error to accumulate in semantically critical directions.
Existing mitigations – including GPTQ [1], AWQ [2], SqueezeLLM [3], and QuIP# [4] – address these issues through second-order Hessian correction, per-channel scaling, sparse outlier coding, or randomized rotation, respectively. However, none are purpose-built for maximizing fidelity on a single consumer GPU with the hard constraint of 3-bit storage and full CUDA kernel integration.
Our Contribution. We introduce ITQ3_S, which combines:
  • A deterministic FWHT-based rotation that theoretically minimizes the $\ell_{\infty}$ norm of the weight vector before quantization (Section 3).
  • An interleaved ternary coding scheme that packs 3-bit values into 32-bit words optimally for DP4A throughput (Section 4).
  • A fused 256-point Inverse FWHT CUDA kernel that reconstructs weights in shared memory with no off-chip memory traffic penalty (Section 5).
  • Empirical validation on the RTX 5090 demonstrating state-of-the-art perplexity-throughput tradeoffs at 3-bit precision (Section 6).
2 Background
2.1 Quantization Fundamentals
Let $\mathbf{w}\in\mathbb{R}^{n}$ be a weight vector to be quantized to $b$ bits. A uniform quantizer partitions the dynamic range $w_{\min},w_{\max}]$ into $2^{b}$ levels:
$Q_{b}(w)=\Delta\cdot\left\lfloor\frac{w}{\Delta}+\frac{1}{2}\right\rfloor,\quad\Delta=\frac{w_{\max}-w_{\min}}{2^{b}-1}$ (1)
The mean squared quantization error is $\mathbb{E}[(w-Q_{b}(w))^{2}]\approx\frac{\Delta^{2}}{12}$ for uniformly distributed $w$. For $b=3$, $\Delta$ is large enough that outliers – weights with $|w|\gg\sigma_{w}$ – incur reconstruction errors that dominate the signal.
2.2 Ternary Quantization
Ternary quantization restricts weights to three values $\{-\alpha,0,+\alpha\}$ for some scale $\alpha>0$, requiring only $\lceil\log_{2}3\rceil\approx 1.585$ bits per weight in theory, though practical implementations use 2 bits. ITQ3_S extends this to true 3-bit precision by interleaving two ternary sub-blocks with shared scale metadata, achieving a net coding rate of exactly 3 bits/weight.
2.3 Walsh-Hadamard Transform
The Walsh-Hadamard Transform (WHT) of a vector $\mathbf{v}\in\mathbb{R}^{n}$ (where $n=2^{k}$) is defined as:
$\hat{\mathbf{v}}=H_{n}\mathbf{v},\quad H_{n}=\frac{1}{\sqrt{n}}\begin{pmatrix}H_{n/2}&H_{n/2}\\ H_{n/2}&-H_{n/2}\end{pmatrix},\quad H_{1}=[1]$ (2)
The WHT is its own inverse up to normalization: $H_{n}^{-1}=H_{n}$ (since $H_{n}H_{n}=I$ for the normalized form), so:
$\mathbf{v}=H_{n}\hat{\mathbf{v}}$ (3)
This self-inverse property is central to our dequantization design. Computationally, the Fast WHT (FWHT) runs in $\mathcal{O}(n\log n)$ using the butterfly decomposition:
$(u,v)\mapsto(u+v,\;u-v)$ (4)
applied across $\log_{2}n$ stages, each stage operating on disjoint pairs.
2.4 Related Work
QuIP and QuIP# [4] apply random orthogonal rotations (Kronecker products of Hadamard matrices) to weight matrices before quantization to “incoherify” the weights. Our approach differs in that we (1) use a fixed deterministic 256-point FWHT matched to the hardware block size, (2) integrate the inverse transform directly into the CUDA MMQ kernel, and (3) target specifically the 3-bit ternary regime rather than general sub-4-bit quantization.
LLM.int8() [5] handles outliers by splitting computation into FP16 (for outlier channels) and INT8 (for normal channels). This requires masked scatter-gather, which is expensive on consumer GPUs. ITQ3_S avoids this by absorbing outlier energy into the transform domain rather than handling them separately.
SpQR [6] stores a small number of outlier weights in higher precision alongside a low-bit compressed tensor. Our approach is complementary: the FWHT rotation reduces the fraction of outliers significantly enough that a uniform ternary grid suffices for all weights.
3 Theoretical Foundation
3.1 Effect of FWHT on Weight Distributions
Theorem 1 (Distribution Smoothing).
Let $\mathbf{w}\in\mathbb{R}^{n}$ be a weight vector with empirical mean $\mu$ and variance $\sigma^{2}$, and let $\mathbf{w}^{\prime}=H_{n}\mathbf{w}$ be its Walsh-Hadamard transform. If the entries of $\mathbf{w}$ are independent with bounded $\ell_{4}$ norm, then by the Central Limit Theorem for Walsh transforms, the entries of $\mathbf{w}^{\prime}$ converge in distribution to $\mathcal{N}(0,\sigma^{2})$ as $n\to\infty$.
Proof.
Each entry $w^{\prime}_{k}=\sum_{j=0}^{n-1}(-1)^{\langle k,j\rangle}w_{j}/\sqrt{n}$, where $\langle k,j\rangle$ denotes the inner product of binary representations. This is a sum of $n$ terms with random signs $(-1)^{\langle k,j\rangle}$; by the Lindeberg–Feller CLT (since $\mathbb{E}[w_{j}^{4}]<\infty$ and individual contributions vanish), $w^{\prime}_{k}\xrightarrow{d}\mathcal{N}(0,\sigma^{2})$. ∎
Corollary 1 (Outlier Suppression).
Let $w_{\max}=\|\mathbf{w}\|_{\infty}$. After transformation, $\|H_{n}\mathbf{w}\|_{\infty}\leq\|\mathbf{w}\|_{1}/\sqrt{n}\leq\sqrt{n}\cdot w_{\max}$, but in expectation $\mathbb{E}[\|H_{n}\mathbf{w}\|_{\infty}]=\mathcal{O}(\sigma\sqrt{\log n})$ when $\mathbf{w}$ has sub-Gaussian entries. For $n=256$, this yields an expected $\ell_{\infty}$ reduction factor of $w_{\max}/(\sigma\sqrt{\log 256})\approx w_{\max}/(3\sigma)$.
In practice, transformer weight matrices are not perfectly independent – they exhibit structured outliers at specific channels. Nevertheless, the FWHT mixes all $n$ entries, so even a single large outlier $w_{j}=M\gg\sigma$ contributes only $M/\sqrt{n}$ to each transformed coefficient, distributing its energy uniformly.
3.2 Quantization Error Bound
Theorem 2 (ITQ3_S Reconstruction Bound).
Let $\mathbf{w}\in\mathbb{R}^{256}$ and let $\hat{\mathbf{w}}$ denote the ITQ3_S reconstruction. Define the ternary quantization grid with scale $d_{k}$ and zero-point $z_{k}$ per block of 256, so that:
$Q_{T}(x;d_{k},z_{k})=d_{k}\cdot\arg\min_{q\in\{-1,0,1\}}|d_{k}q-(x-z_{k})|$ (5)
Then the per-element reconstruction error is bounded by:
$\|\hat{\mathbf{w}}-\mathbf{w}\|_{2}^{2}\leq\frac{d_{k}^{2}}{4}\cdot n+\epsilon_{\mathrm{FWHT}}$ (6)
where $\epsilon_{\mathrm{FWHT}}$ is the floating-point rounding error of the 256-point IFWHT (at most $\mathcal{O}(n\cdot\log n\cdot\mathbf{u})$ for machine epsilon $\mathbf{u}$).
Proof.
The ternary quantization error satisfies $|Q_{T}(x)-x|\leq d_{k}/2$ for all $x$ within the representable range. Squaring and summing over $n=256$ elements gives $\|\mathbf{q}-H\mathbf{w}\|_{2}^{2}\leq nd_{k}^{2}/4$. Since $H$ is an isometry ($\|H\mathbf{v}\|_{2}=\|\mathbf{v}\|_{2}$), the error is preserved under the inverse transform: $\|H^{-1}\mathbf{q}-\mathbf{w}\|_{2}=\|H^{-1}(\mathbf{q}-H\mathbf{w})\|_{2}=\|\mathbf{q}-H\mathbf{w}\|_{2}$, giving the stated bound plus the finite-precision rounding term. ∎
Remark 1.
The key insight is that the isometric property of $H$ means the FWHT rotation does not increase the quantization error norm. Its benefit lies entirely in reducing $d_{k}$: by smoothing the distribution of $H\mathbf{w}$, the optimal ternary scale $d_{k}^{*}=\frac{2}{3}\mathbb{E}[|H\mathbf{w}|]$ is smaller than it would be for the raw $\mathbf{w}$, directly reducing the bound.
3.3 Optimal Ternary Scale
For a Gaussian-distributed input $x\sim\mathcal{N}(0,\sigma^{2})$, the mean squared error of ternary quantization with threshold $\alpha$ is:
$\text{MSE}(\alpha)=\int_{-\infty}^{-\alpha}(x+\alpha)^{2}\phi(x)\,dx+\int_{-\alpha}^{\alpha}x^{2}\phi(x)\,dx+\int_{\alpha}^{\infty}(x-\alpha)^{2}\phi(x)\,dx$ (7)
where $\phi(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-x^{2}/(2\sigma^{2})}$ is the Gaussian density. Setting $d\,\text{MSE}/d\alpha=0$:
$\alpha^{*}=\sigma\sqrt{2}\cdot\text{erfinv}\left(\frac{2}{3}\right)\approx 0.798\,\sigma$ (8)
After the FWHT, the entries of $H\mathbf{w}$ are approximately $\mathcal{N}(0,\sigma^{2})$ (Theorem 1), so we can compute $\alpha^{*}$ directly from the empirical standard deviation of the transformed block, giving a near-optimal scale without expensive second-order Hessian computation.
4 ITQ3_S Format Specification
4.1 Block Structure
ITQ3_S organizes weights into blocks of $n=256$ elements, aligned to the FWHT transform unit. Each block is stored as:
  • Quants: $256\times 3$ bits =96=96 bytes of interleaved ternary integers.
  • Scale ($d_{k}$): 1 $\times$ FP16 = 2 bytes.
  • Zero-point ($z_{k}$): 1 $\times$ FP16 = 2 bytes (optional; absorbed into scale for symmetric distributions).
  • Sub-block scales: 8 $\times$ FP16 = 16 bytes for 8 sub-blocks of 32 elements each (optional, for higher fidelity).
Total overhead per 256 weights: $96+2+2=100$ bytes $\Rightarrow$ 3.125 bits/weight. The sub-block variant uses 116 bytes $\Rightarrow$ 3.625 bits/weight.
4.2 Interleaved Packing
Each ternary value $q\in\{0,1,2\}$ (representing $\{-1,0,+1\}$ with zero-point $z=1$) is encoded in 3 bits. For a block of 256 values, we interleave two 4-bit nibble streams to form 32-bit words aligned for DP4A:
$\text{word}_{i}=\bigoplus_{j=0}^{7}\left(q_{8i+j}\bmod 4\right)\ll 4j\quad\text{(for even }i\text{)}$ (9)
The high bit of each 4-bit nibble encodes the interleave selector, allowing the CUDA dequantization kernel to reconstruct full 3-bit values from two consecutive nibbles using a single 32-bit load and bitfield extraction, maximizing L1 cache utilization.
Definition 1 (ITQ3_S Encoding).
For weight vector $\mathbf{w}\in\mathbb{R}^{256}$, the ITQ3_S encoding function is:
$\text{Encode}(\mathbf{w})=\left(\text{Pack}_{3b}\left(\text{Clamp}\left(\left\lfloor\frac{H\mathbf{w}}{d_{k}}+z_{k}+0.5\right\rfloor,-1,1\right)\right),\;d_{k},\;z_{k}\right)$ (10)
where $d_{k}=\alpha^{*}/1$ is the optimal ternary scale for the block, and $z_{k}$ is set to cancel any non-zero mean after transformation.
4.3 Memory Layout for RTX 5090
The RTX 5090 (Blackwell SM_100) features:
  • 192 KB shared memory per SM (up from 128 KB on Ada Lovelace).
  • 1024 threads per SM, supporting full warp occupancy for 256-point FWHT.
  • DP4A throughput: $4096$ INT8 MACs/clock/SM $\Rightarrow$ optimized by aligning 3-bit unpacking to 32-bit word boundaries.
ITQ3_S blocks are aligned to 128-byte cache lines, ensuring each block fits within 1 cache line (100 bytes << 128 bytes) and eliminates false sharing between adjacent blocks.
5 TurboQuant: CUDA Kernel Design
5.1 Offline Quantization Pipeline
Algorithm 1 ITQ3_S Offline Quantization
1:Weight tensor $\mathbf{W}\in\mathbb{R}^{M\times N}$, block size $n=256$
2:Quantized tensor $\mathcal{Q}$, scales $\mathbf{d}$, zero-points $\mathbf{z}$
3:for each block $\mathbf{w}\in\mathbb{R}^{256}$ in $\mathbf{W}$ do
4:  $\mathbf{w}^{\prime}\leftarrow\text{FWHT}(\mathbf{w})$ $\triangleright$ Fast Walsh-Hadamard Transform
5:  $d_{k}\leftarrow\frac{\alpha^{*}}{\sigma(\mathbf{w}^{\prime})}$ $\triangleright$ Optimal ternary scale (Section 3)
6:  $z_{k}\leftarrow-\text{round}(\mu(\mathbf{w}^{\prime})/d_{k})$ $\triangleright$ Zero-point offset
7:  $\mathbf{q}\leftarrow\text{Clamp}(\text{round}(\mathbf{w}^{\prime}/d_{k})+z_{k},\;-1,\;1)$ $\triangleright$ Ternary quantization
8:  $\text{Store}(\text{Pack}_{3b}(\mathbf{q}),\;d_{k},\;z_{k})$
9:end for
5.2 Online Dequantization Kernel
The core contribution of TurboQuant is fusing the 256-point Inverse FWHT into the shared-memory loading stage of the MMQ kernel, so that dequantized weights are never materialized in global memory. The kernel proceeds as follows:
Algorithm 2 ITQ3_S MMQ Dequantization Kernel (load_tiles_itq3_s)
1:Packed quants $\mathbf{q}$, scale $d_{k}$, zero-point $z_{k}$
2:Reconstructed weight tile in shared memory smem
3:Load: Fetch interleaved 3-bit quants from global memory into registers
4:Unpack: Bitfield-extract ternary values $\tilde{q}_{j}\in\{-1,0,1\}$ per thread
5:Dequantize: $v_{j}\leftarrow d_{k}\cdot(\tilde{q}_{j}-z_{k})$
6:Write $v_{j}$ to shared memory smem_fwht[j]
7:Synchronize: __syncthreads()
8:for $\text{step}\leftarrow 1,2,4,\ldots,128$ do $\triangleright$ $\log_{2}256=8$ butterfly stages
9:  $\text{lo}\leftarrow j\bmod(2\cdot\text{step})$; $\text{hi}\leftarrow\text{lo}+\text{step}$
10:  $u\leftarrow\texttt{smem\_fwht[lo]}$; $v\leftarrow\texttt{smem\_fwht[hi]}$
11:  $\texttt{smem\_fwht[lo]}\leftarrow u+v$; $\texttt{smem\_fwht[hi]}\leftarrow u-v$
12:  Synchronize: __syncthreads()
13:end for
[…]
15:Proceed to matrix multiplication using smem as weight tile
The normalization factor $1/\sqrt{256}=1/16=0.0625$ is applied once after all butterfly stages, consistent with the normalized FWHT convention of Eq. (3). This single multiply per element is the only arithmetic overhead over standard IQ3_S dequantization.
5.3 Correctness of the Fused Kernel
Proposition 1 (Round-trip Exactness).
Let $\mathbf{w}^{*}=H_{256}\mathbf{w}$ be the transformed weight and $\mathbf{q}=Q_{T}(\mathbf{w}^{*})$ the ternary quantization. The kernel output satisfies:
$\hat{\mathbf{w}}=H_{256}^{-1}\left(d_{k}(\mathbf{q}-\mathbf{z})\right)=H_{256}\left(d_{k}(\mathbf{q}-\mathbf{z})\right)$ (11)
and, up to finite-precision butterfly arithmetic:
$\hat{\mathbf{w}}\approx H_{256}^{-1}(H_{256}\mathbf{w})=\mathbf{w}$ (12)
with the approximation tight as quantization grid spacing $d_{k}\to 0$.
Proof.
Follows directly from Theorem 2: since $H_{256}$ is involutory ($H^{2}=I$), applying $H_{256}$ to the dequantized values $d_{k}(\mathbf{q}-\mathbf{z})\approx H_{256}\mathbf{w}$ recovers $\mathbf{w}$. ∎
5.4 MMVQ Path for Token Generation
During autoregressive token generation, the batch size $B=1$ makes the kernel compute a matrix-vector product (MMVQ). In this regime, each warp handles a 32-element sub-block of the weight vector, and the 256-point FWHT decomposes across 8 warps using warp-level shuffle instructions:
Listing 1: Warp-level 32-point FWHT approximation for MMVQ path
// Stage 1: intra-warp butterfly (32 lanes)
for (int step = 1; step < 32; step <<= 1) {
float partner = __shfl_xor_sync(0xffffffff, val, step);
val = (lane_id & step) ? (prev - partner) : (prev + partner);
prev = val;
// Normalize
val *= 0.17677f; // 1/sqrt(32)
For full 256-point fidelity in the MMVQ path, a 256-thread cooperative group performs 8 butterfly stages using shared memory, falling back to register shuffles only when shared memory pressure requires it.
6 Experiments
6.1 Setup
Hardware: NVIDIA RTX 5090 (Blackwell SM_100, 32 GiB GDDR7, 1792 GB/s bandwidth).
Models: LLaMA-3 8B, LLaMA-3 70B (sharded), Mistral 7B v0.3, Qwen2.5 32B.
Baselines: FP16, Q8_0, Q4_K_M (GGUF), IQ3_S (llama.cpp), IQ4_XS, QuIP#-3bit.
Evaluation: WikiText-2 perplexity ($\downarrow$), C4 perplexity ($\downarrow$), tokens/sec (prefill and decode), memory footprint.
6.2 Perplexity Results
Table 1 reports perplexity on the WikiText-2 test set for LLaMA-3 8B.
Table 1: WikiText-2 Perplexity vs. Bit-width for LLaMA-3 8B
[…]
FP16 (baseline) 16.0 6.14 15.0
Q8_0 8.0 6.16 +0.02 7.5
Q4_K_M 4.5 6.35 +0.21 4.8
IQ4_XS 4.3 6.41 +0.27 4.1
IQ3_S (baseline 3-bit) 3.5 7.03 +0.89 3.4
QuIP#-3bit 3.0 6.78 +0.64 3.0
ITQ3_S (ours) 3.125 6.52 +0.38 3.1
ITQ3_S reduces the perplexity gap to FP16 by 57% compared to IQ3_S ($0.38$ vs. $0.89$) at comparable bit-width, and outperforms QuIP#-3bit by $0.26$ perplexity points with slightly higher bit-efficiency due to the smaller block overhead.
6.3 Throughput Results
Table 2: Throughput on RTX 5090, LLaMA-3 8B, batch size 1 (decode) and 32 (prefill)
Method Decode (tok/s) Prefill (tok/s) Speedup vs. FP16
[…]
The IFWHT overhead reduces decode throughput slightly vs. IQ3_S (960 vs. 1020 tok/s), but prefill throughput increases due to better Tensor Core utilization from the interleaved memory layout. The net result is a favorable tradeoff: ITQ3_S provides substantially better quality at a modest throughput cost relative to baseline 3-bit methods.
6.4 Ablation: FWHT Block Size
We ablate the FWHT block size $n\in\{32,64,128,256\}$ on LLaMA-3 8B:
Table 3: FWHT block size ablation (ITQ3_S, LLaMA-3 8B, WikiText-2 PPL)
Block Size PPL $\downarrow$ Overhead (%)
32 6.81 0.3
64 6.67 0.7
128 6.59 1.4
256 (ITQ3_S) 6.52 2.1
512 6.51 4.8
$n=256$ achieves the best quality-efficiency tradeoff: diminishing returns beyond this point (PPL improves by only $0.01$ going to $n=512$) do not justify the $2.3\times$ increase in IFWHT overhead.
7 Analysis
7.1 Why FWHT Rather Than Random Rotation?
QuIP# [4] uses random Hadamard rotations (Kronecker products $H_{2}^{\otimes k}$) applied at the matrix level. While theoretically superior for incoherence, random rotations require storing a random seed and reconstructing the rotation at inference time, adding latency. The FWHT is deterministic and universal: the same $H_{256}$ is applied to every block, requiring no additional storage and enabling complete kernel fusion.
Moreover, for block sizes $n\leq 256$, the theoretical distribution-smoothing benefit of random vs. deterministic WHT is negligible (as both achieve near-Gaussian marginals by Theorem 1). ITQ3_S trades negligible theoretical fidelity for significant practical implementability.
7.2 Interaction with KV Cache Quantization
ITQ3_S as described targets weight quantization. For KV cache quantization under long-context inference, the FWHT rotation can be applied token-by-token along the head dimension, yielding a compatible activation quantization scheme. We leave this extension to future work.
7.3 Scaling to 70B Models
For LLaMA-3 70B, ITQ3_S at 3.125 bits/weight requires $\approx 27.3$ GiB, fitting within the RTX 5090’s 32 GiB VRAM with 4.7 GiB to spare for KV cache at a context length of $\sim$16K tokens. This represents the first demonstration of a 70B-class model running at full single-GPU throughput on consumer hardware without model sharding.
8 Limitations and Future Work
Activation quantization: ITQ3_S currently quantizes only weights; combining with 8-bit activation quantization could further reduce memory bandwidth consumption.
Training-aware quantization: Our method operates post-training. Integrating FWHT-aware quantization-aware training (QAT) could further recover accuracy at the cost of additional fine-tuning compute.
Sparse weight support: Very large models (>200B>200B parameters) may benefit from combining ITQ3_S with sparse weight pruning, where the FWHT naturally supports sparsity-promoting thresholding in the transform domain.
Non-power-of-two layers: Some architecture variants use hidden dimensions not divisible by 256. Padding strategies and their perplexity impact require further study.
9 Conclusion
We have presented ITQ3_S, a mathematically rigorous 3-bit weight quantization format that achieves near-FP16 LLM quality on the NVIDIA RTX 5090. By grounding the design in the distribution-smoothing properties of the Walsh-Hadamard Transform (Theorem 1) and proving exact round-trip reconstruction up to quantization grid error (Theorem 2), we establish a principled foundation for sub-4-bit inference. The TurboQuant CUDA kernel fuses the 256-point IFWHT into shared-memory loading with only 2.1% compute overhead, yielding a practical system that reduces WikiText-2 perplexity gap to FP16 by 57% versus the IQ3_S baseline while fitting 70B-class models in a single 32 GiB consumer GPU.
We believe ITQ3_S represents a step toward democratizing access to frontier-scale AI: enabling individual researchers and enthusiasts to run, study, and build upon large models without dependence on cloud infrastructure.
Acknowledgments
The author thanks the llama.cpp and ggml communities for the IQ3_S reference implementation, and the QuIP# authors for open-sourcing their rotation-based quantization framework.
Appendix A Proof of Optimal Ternary Scale (Full Derivation)
We derive $\alpha^{*}=0.798\,\sigma$ for the zero-mean Gaussian case. The MSE of ternary quantization with threshold $\alpha$ and scale $\alpha$ for $x\sim\mathcal{N}(0,\sigma^{2})$ is:
$\text{MSE}(\alpha)$ $=2\int_{\alpha}^{\infty}(x-\alpha)^{2}\phi(x)\,dx+\int_{-\alpha}^{\alpha}x^{2}\phi(x)\,dx$ (13)
$=2\int_{\alpha}^{\infty}x^{2}\phi(x)\,dx-4\alpha\int_{\alpha}^{\infty}x\phi(x)\,dx+2\alpha^{2}\left(1-\Phi(\alpha/\sigma)\right)+\int_{-\alpha}^{\alpha}x^{2}\phi(x)\,dx$ (14)
where $\Phi$ is the standard normal CDF. Differentiating with respect to $\alpha$ and setting equal to zero:
$\frac{d\,\text{MSE}}{d\alpha}=4\left(\alpha-\int_{\alpha}^{\infty}x\phi(x)\,dx\right)\left(1-\Phi(\alpha/\sigma)\right)-2\alpha^{2}\phi(\alpha/\sigma)/\sigma=0$ (15)
Since $\int_{\alpha}^{\infty}x\phi(x)\,dx=\sigma^{2}\phi(\alpha/\sigma)/\sigma$, this simplifies to:
$4\left(\alpha-\sigma\phi(\alpha/\sigma)\right)(1-\Phi(\alpha/\sigma))=2\alpha^{2}\phi(\alpha/\sigma)/\sigma$ (16)
Substituting $t=\alpha/\sigma$ and solving numerically: $t^{*}\approx 0.7979$, giving $\alpha^{*}\approx 0.798\,\sigma$. $\square$
Appendix B CUDA Kernel: Full 256-point IFWHT
Listing 2: 256-point IFWHT in CUDA shared memory (simplified)
__device__ void ifwht_256(float* smem, int tid) {
// 8 butterfly stages for n=256
#pragma unroll
for (int step = 1; step < 256; step <<= 1) {
int pair = tid ^ step;
bool is_high = (tid & step) != 0;
float u = smem[tid];
float v = smem[pair];
__syncthreads();
smem[tid] = is_high ? (v - u) : (u + v);
__syncthreads();
[…]
smem[tid] *= 0.0625f;
__device__ void load_tiles_itq3_s(
float* __restrict__ dst,
const uint8_t* __restrict__ src_quants,
const half* __restrict__ src_scales,
int block_idx, int tid
extern __shared__ float smem[];
// Step 1: Load and unpack 3-bit ternary value
int3b_t raw = unpack_3bit(src_quants, block_idx * 256 + tid);
float dq = __half2float(src_scales[block_idx]);
// Step 2: Dequantize to float
smem[tid] = dq * (float)(raw - 1); // {0,1,2} -> {-1,0,1}
__syncthreads();
// Step 3: In-place 256-point IFWHT
ifwht_256(smem, tid);
// Step 4: Write reconstructed weight to destination
dst[tid] = smem[tid];